OpenStudy (anonymous):

When you solve an equation, how do you know when to add and when to subtract? You may include an example if you need to.

OpenStudy (anonymous):

x + 2 =1 to get the value of x you should subtract 2 from 1 x = 1-2 x=-1 okay??